In 11 years, Rose will be 5 times as old as she was 9 years ago. How old is she now?
Answer by stanbon(75887)

In 11 years, Rose will be 5 times as old as she was 9 years ago. How old is she now?
----
Equation:
r + 11 = 5(r-9)
r + 11 = 5r -45
4r = 56
r = 14 (Rose's age now)
--------------------
Cheers,
Stan H.